# The Face of Populism: Examining Differences in Facial Emotional Expressions of Political Leaders Using Machine Learning

### v4

@@ Abstract
Populist rhetoric employed on online media is characterized as deeply impassioned and often imbued with strong emotions. The aim of this paper is to empirically investigate the differences in affective nonverbal communication of political leaders. We use a deep-learning approach to process a sample of 220 YouTube videos of political leaders from 15 different countries, analyze their facial expressions of emotion and then examine differences in average emotion scores representing the relative presence of 6 emotional states (anger, disgust, fear, happiness, sadness, and surprise) and a neutral expression for each frame of the YouTube video. Based on a sample of manually coded images, we find that this deep-learning approach has 53-60% agreement with human labels. We observe statistically significant differences in the average score of negative emotions between groups of leaders with varying degrees of populist rhetoric.
@@ 1 Introduction
In the growing body of literature on populism, the category of emotion has most frequently been utilized in discrete ways to describe its emergence and tenacity (Demertzis 2006)—by reference to the disillusioned silent majority, heightened emotions surrounding political leaders, and emotionally charged divisions between the people and the elite (Moffitt & Tormey 2014, Mudde 2017, Norris & Inglehart 2019). Compared to pluralist parties, who are commonly portrayed as more affectively neutral, populist rhetoric is most often characterized as deeply impassioned and imbued with strong (negative) emotions (Bonansinga 2020a, Salmela & von Scheve 2018). Although the nature of the relationship between emotionality and populist discourse has rarely been subject to systematic empirical scrutiny (Widmann 2021), this link has become one of the main focal points of research on political communication in recent years. Comprehensive research on the matter is thus crucial for gaining insight into the extent to which populist narratives and performances are rooted in and shaped by emotion, and by extension, how emotion may be utilized as a tool for building and maintaining political support (Nguyen 2019).
@@
In the past, studies of populist communication have generally focused on its verbal elements, by analyzing the affective content of the words used in oral or written addresses of political leaders (Bobba 2019, Breeze 2019, 2020, Cochrane et al. 2022, Jacobs et al. 2020, Martella & Bracciale 2022, Widmann 2021). While advantageous in its own way, this type of analysis can pose a challenge to researchers primarily due to language barriers and the fact that written text usually contains only a small amount of words that have an inherent affective connotation (Widmann 2021). At the same time, currently available sentiment analysis tools are predominantly based on and developed for the English language, while the output for other languages is often of lower quality (Mohammad 2016). These circumstances therefore limit the possibility of result generalization and cross-cultural comparison.
@@
Visual analyses, although insufficiently represented in the given field of research, may provide further insight into the characteristics of affective populist communication as they are not hindered by the same limitations as text-based studies (Moffitt 2022b) and allow for analysis of the visual rhetoric constructed and utilized by political leaders (Mendonça & Caetano 2021). Within the scope of visual research on political emotion display, two main schemes of analysis have shown to be especially valuable: the ethological approach (Masters et al. 1986) and the Facial Action Coding System (FACS) (Ekman & Friesen 1976).
@@
Studies within the former largely focus on the analysis of political candidates’ non-verbal representations in news coverage, their displays during political campaigns and debates, as well as the impact of different forms of representation on viewers and success in political elections (Bucy et al. 2020, Bucy & Grabe 2007, 2008, Grabe & Bucy 2009, Sullivan & Masters 1988). The Facial Action Coding System, on the other hand, allows for a more in-depth analysis of the (micro)expressive behavior of political leaders during speeches and public performances. Due to its complexity, Ekman’s classification system has been applied in a limited number of studies (Stewart et al. 2009, 2015, Stewart & Ford Dowe 2013) but as the process of visual analysis gradually becomes automated (Joo et al. 2019) this research method is becoming more accessible to researchers from various fields of interest (Ash & Caliskan 2022, Bossetta & Schmøkel 2023).
@@
Our study expands on the existing framework of FACS research on political emotion display and applies a computer vision approach for detecting emotions from the facial expressions of political leaders. Because it is relatively new, this approach faces various methodological challenges which we aim to address by following the recommendations outlined by Barrett et al. (2019), such as using big data techniques and automated detection in naturalistic settings, integrating traditional methods with machine learning, and sampling visual data across diverse contexts and cultures.
@@
Since the introduction of the FER dataset (Goodfellow et al. 2013), different machine learning algorithms have been developed for solving the challenge of facial expression recognition, which essentially requires performing two tasks: (1) detection of facial expression, and (2) emotion recognition from the detected facial expressions. We develop a reproducible and principled computational workflow in Python for the analysis of video content of political leaders’ public performances. Our work relies on the existing Python library fer (Shenk et al. 2021) which bundles a pre-trained convolutional neural network (CNN) Keras models for face detection (Zhang et al. 2016a, based on) and emotion recognition (Arriaga et al. 2017, 2020).
@@ 2 Populism as performance
Defining populism represents a challenging task for researchers as there is barely any consensus on what such a definition should (or should not) include, and the concept itself becomes increasingly muddled as the term is used to describe a growing number of phenomena (Gidron & Bonikowski 2013). Most conceptualizations of populism can, however, be classified (Rovira Kaltwasser et al. 2017) into one of three main approaches: (1) the ideational approach (Mudde 2004), (2) the discursive approach (Norris & Inglehart 2019, Ostiguy 2017) and (3) the political-strategic approach (Weyland 2017). For the purposes of this study, we adopt the discursive approach, which conceives populism primarily as a rhetorical style based on first-order principles pertaining to who should rule—the people, as “the only legitimate source of political and moral authority in a democracy”—but which says very little about the how (Norris & Inglehart 2019). Whereas the ideational approach claims populism is a “thin-centered” ideology (Mudde 2004) and emphasizes the interpretation of politics as “a Manichean struggle between a reified will of the people and a conspiring elite” (Hawkins & Kaltwasser 2018) (Hawkins and Rovira Kaltwasser, 2018), the discursive approach highlights the ways in which this interpretation is framed and communicated as an anti-establishment discourse (Aslanidis 2017). We find this to be the most comprehensible understanding of populism as it considers the operationalized elements of the populist communication logic (Engesser et al. 2017) and examines the rhetorical components implicit in the ideological definition (Aslanidis 2016).
@@
From this perspective, the populist rhetoric has widely been characterized by the use of “dramatization, polarization, moralization, directness, ordinariness, colloquial and vulgar language” (Engesser et al. 2017), and low-order appeals within which leaders “use a language that includes slang or folksy expressions and metaphors, are more demonstrative in their bodily or facial expressions as well as in their demeanor, and display more raw, culturally popular tastes” (Ostiguy 2017). Because it represents a form of symbolic resistance to the more formalized and composed self-representation of mainstream pluralist leaders, i.e., the “establishment” and “political elite”, the populist performance  [^1] This does not imply that, on average, there is a tendency for Japanese leaders to be more neutral. In fact, in Japan, we observe the same tendency as seen in the entire sample — the speeches of Shinzo Abe (LDP) from the populist group have the highest average score of negative emotions (0.58), while Natsuo Yamaguchi (NKP) from the moderately pluralist group has the lowest average score (0.30). Conversely, the average scores of neutrality show the opposite trend. These results are not extreme compared to the sample distributions shown in 2.is subversive to the normative communication structures of political dialogue. As such, it appeals to “common sense” (Bucy et al. 2020) and claims to give a voice to “the ordinary people” (Inglehart & Norris 2016).
@@
This mode of communication has shown to be highly effective as a means of garnering media attention and cultivating a dedicated base of supporters—especially in the new media environment—and populist leaders seem to be overall more adept at capitalizing on it than other, non-populist political figures (Bobba 2019, Davidson & Enos 2022). Since digital media hinges on the attention economy (Wells et al. 2020) and thus tends to favor “dark participation” (Larsson 2022), populist leaders and political parties may fare well online due to the transgressive nature of their self-representation and the simplification, emotionalization, and negativity inherent in populist messaging (Bucy et al. 2020, Engesser et al. 2017). Social media is a particularly convenient tool for the proliferation of populist messages and imagery as it allows political actors to circumvent the mainstream media, reach large audiences, and communicate and interact directly with ‘the people’ (Lalancette & Raynauld 2019, Moffitt 2022a). At the same time, online platforms allow for an instant audience response, as the “feedback loop between producer and consumer is much more intense and rapid than with television and radio (Finlayson 2022). This form of ‘mediatized populism’ (Mazzoleni 2014) allows for aspecific type of political performance, one which is largely unmitigated and relies heavily on portraying political leaders favorably to (online) audiences by framing them as expressive, dominant, and disruptive figures who are in tune with “the will of the people”.
@@
In recent years, much attention has been paid to the strategies of (affective) populist communication online, especially through analyses of verbal content shared by populist leaders and parties on Twitter and Facebook (e.g. Bracciale & Martella 2017, Gründl 2022, Jacobs et al. 2020, Martella & Bracciale 2022, Widmann 2021). Given the rise of visual communication in politics and the rapid developments of Web 2.0 practices (Grabe & Bucy 2009, Larsson 2022, Messaris 2019), our primary aim was to expand the scope of research on the affective self-representation of populist leaders online by examining the characteristics of their non-verbal emotional communication. The facial expressions of political leaders are a particularly valuable area of research in this context, since digital media provide a breadth of material for analysis—well-lit, high-quality close-up shots of political leaders are now more prevalent than ever. We used this abundance of available data sources to our advantage and processed 220 YouTube videos uploaded to official accounts of political parties and their leaders, amounting to a total of 77 hours of video material. Our analysis examines potential differences in facial expressions between populist and non-populist political leaders during public performances, and more specifically, the characteristics of their negative emotional expressions and neutrality in emotional stance (i.e., absence of emotional expression).
@@ 3 Data & Method
To create our sample, we first had to differentiate and categorize political parties (and, by extension, their leading figures) as either populist or pluralist. For this purpose, we used data from the Global Party Survey (GPS) (Norris 2020), an expert survey from 2019 whichoffers an overview of key ideological and issue positions for 1043 parties from 163 countries worldwide. While expert surveys certainly have their limitations (Wiesehomeier 2018) and may not be able to fully capture the nuances of populist rhetoric, this database remains the most comprehensive comparative overview of political parties globally. Among other data, such as ideological framework and party size, the survey also offers an estimate of the degree to which existing parties employ populist rhetoric. In the questionnaire design, populist rhetoric is defined as “language which typically challenges the legitimacy of established political institutions and emphasizes that the will of the people should prevail” (Norris 2020), whereas pluralist rhetoric “rejects these ideas, believing that elected leaders should govern, constrained by minority rights, bargaining and compromise, as well as checks and balances on executive power” (Norris 2020). The populist rhetoric scale in (V8 in GPS dataset) is based on expert assessments of rhetorical style given these parameters. The variable Type_Populism categorizes this scale into four ordinal groups: strongly pluralist, moderately pluralist, moderately populist, and strongly populist. We also use the “Type_Values” variable in supplementary analysis to examine ideological differences in expression of emotions (see Supplement section E – Ideological differences).
@@
From the GPS dataset, fifteen countries with competitive popular elections were chosen: Argentina, Australia, Brazil, Croatia, France, Hungary, India, Italy, Japan, the Netherlands, Serbia, South Africa, Turkey, the United Kingdom, and the United States. This sample, although limited, provides somewhat broad coverage of various cultures and political systems globally. Three parties with varying degrees of populist rhetoric—strongly populist, strongly pluralist, and either moderately populist or moderately pluralist—were selected from each country. We primarily focused on minor and major political parties as defined by the GPS variable categorizing parties based on their share of the vote in contests for the lower house of the national parliament. Fringe parties (less than 3% of the vote) were generally excluded from the sample since their audience reach and influence is considerably less than that of their larger counterparts competing at the same scale in their respective countries. Likewise, due to this criteria, only two parties from the US were included in the sample. Table LABEL:tab:si-lead from the Supplementary Information contains information about the selected parties and their leaders.
@@
We then identified one leader per party who has held a prominent political position over the past decade (up to 2019, covered by the GPS), and found their official YouTube channel or their party’s official YouTube channel. From these channels, five videos were selected for each leader based on the following criteria:
@@
    View count. Videos that reached a larger audience were given priority over videos that did not perform comparatively as well on the platform. View count was used as a filtering mechanism to single out videos with large and potentially diverse audiences, because the most viewed videos on the official YouTube channels of major political parties usually refer to topics of wider public interest. This is partially due to the “popularity bias” on YouTube, meaning it has a tendency to recommend political content that aligns with the interests of the majority (Heuer et al. 2021).
@@
We processed each video in the sample by extracting 300 frames uniformly distributed over the length of the video, which were the basis for our main analysis. To improve the robustness of the overall approach, we also included an alternative frame processing approach in the supplementary material, based on processing every 50th frame of the video. Given that the majority of the YouTube videos have 24–30 frames per second, this is equivalent to selecting a frame every 1.67–2.00 seconds, which is a reasonable time window if we expect to capture potential changes in facial expression. The average number of frames per video extracted using this approach was 273.03, which makes the resulting time-series of emotion scores comparable in length with the first approach.
@@
For each selected frame of every video in the sample the following procedure is applied. First, face detection is applied to the extract frames using a multi-task cascaded framework based on convolutional neural networks (CNN) (Zhang et al. 2016b). This framework is suitable for detecting faces in videos because it performs simultaneous face detection and alignment, which is important as we cannot expect that political leaders will be looking straight at the camera in each extracted video frame. After face detection was applied to all frames, there were videos in which multiple faces were detected throughout the frame. These videos were manually checked to confirm which of the detected faces corresponds to the selected political leader.
@@
The second step is emotion detection within the verified detected face box. We relied on implementing the mini-Xception architecture of CNN (Arriaga et al. 2017) aimed at detecting 6 basic facial expressions of emotion, as well as a neutral expression. Similar to the face detection approach outlined above, this implementation is aimed at real-time emotion detection from video input for robotic applications. This model achieves 66% accuracy (see Arriaga et al. 2017, p. 3 for more details) on the FER2013 dataset (Goodfellow et al. 2013) which is a standard benchmark for emotion detection. The dataset consists of 30,000 small images (48x48 pixels) showing diverse facial expressions of different emotions.  [^2] FER2013 was constructed using Google Images API queries for emotion-related keywords like “blissful” and “enraged”, combined with queries introducing diversity in gender, age and ethnicity represented in the images. Images were then labeled by human coders (Goodfellow et al. 2013, pp. 3-4). Unfortunately, no information is provided by the authors of the dataset regarding the number of human coders nor their background, or the labeling methodology applied to extract training labels for every image included in the FER2013 dataset. Benchmarks based on the FER2013 dataset should be taken with caution, especially in social science applications, because the dataset construction and validation documentation lacks crucial information about inter-coder reliability, coding procedures, and potential biases in the image selection process. Each emotional expression has approximately 6000 images, except disgust, which is only present in 600 images. According to Goodfellow et al. (2013), human classification accuracy is in the range of $65\%\pm 5\%$, but no additional information about the sample size, number of coders, or detection procedure is given.
@@
In order to provide additional information and to validate the output of the machine learning approach, we conducted a small image annotation study where 5 coders annotated a sample of 514 randomly selected frames extracted from 220 videos. Coders received basic training in emotion recognition and, due to limited resources, answered only two questions about every image: (1) Can you clearly see the face of a politician in this image? (Yes/No), (2) Is this face expressing positive emotion, negative emotion, or is it a neutral expression? (“Positive emotion”, “Negative emotion”, “Neutral expression”). The majority vote label was set as the “human-coded label” and compared with the output of the machine learning algorithm, whose label is defined based on the highest score out of the sum of positive emotions, negative emotions, or the neutrality score. A detailed description and the results of the data annotation study can be found in Supplement F.
@@
Figure 1: Example frames showing a high detected score of negative emotions (top: angry 0.78, disgust 0.00, fear 0.09, happy 0.00, sad 0.07, surprise 0.00, neutral 0.04) and a high detected score of neutral expression (bottom: angry 0.07, disgust 0.00, fear 0.04, happy 0.00, sad 0.02, surprise 0.08, neutral 0.79).

### v5

@@ Abstract
Populist rhetoric employed in online media is characterized as deeply impassioned and most often imbued with strong emotions. This paper investigates the differences in affective non-verbal communication of political leaders. We use a deep-learning approach to process a sample of 220 YouTube videos depicting political leaders from 15 different countries, analyze their facial expressions of emotion, and then examine differences in average emotion scores representing the relative presence of 6 emotional states (anger, disgust, fear, happiness, sadness, and surprise) and a neutral expression for each frame of the processed YouTube video. Based on a sample of manually coded images, we find that this machine learning approach has 53-60% agreement with human annotation. We observe statistically significant differences in the average score of expressed negative emotions between groups of leaders with varying degrees of populist rhetoric. Overall, our contribution provides insight into the characteristics of non-verbal emotional expression among political leaders, as well as an open-source workflow for further computational studies of their affective communication.
keywords:
populism, facial expressions, emotion recognition, machine learning, political communication
Please refer to and cite the published version:
Major, S., Tomašević, A. The face of populism: examining differences in facial emotional expressions of political leaders using machine learning. J Comput Soc Sc 8, 62 (2025). https://doi.org/10.1007/s42001-025-00392-w
Abstract
@@ 1 Introduction
In the growing body of literature on populism, the category of emotion has most frequently been utilized in discrete ways to describe its emergence and tenacity [1]—by reference to the disillusioned silent majority, heightened emotions surrounding political leaders, and emotionally charged divisions between the people and the elite [2, 3, 4]. Compared to pluralist parties, who are commonly portrayed as more affectively neutral, populist rhetoric is most often characterized as deeply impassioned and imbued with strong (negative) emotions [5, 6]. Although the nature of the relationship between emotionality and populist discourse has rarely been subject to systematic empirical scrutiny [7], this link has become one of the main focal points of research on political communication in recent years. Comprehensive research on the matter is thus crucial for gaining insight into the extent to which populist narratives and performances are rooted in and shaped by emotion, and by extension, how emotion may be utilized as a tool for building and maintaining political support [8].
@@
In the past, studies of populist communication have generally focused on its verbal elements, by analyzing the affective content of the words used in oral or written addresses of political leaders [9, 10, 11, 12, 13, 14, 7]. While advantageous in its own way, this type of analysis can pose a challenge to researchers primarily due to language barriers and the fact that written text usually contains only a small amount of words that have an inherent affective connotation [7]. At the same time, currently available sentiment analysis tools are predominantly based on and developed for the English language, while the output for other languages is often of lower quality [15]. These circumstances therefore limit the possibility of result generalization and cross-cultural comparison.
@@
Visual analyses, although insufficiently represented in the given field of research, may provide further insight into the characteristics of affective populist communication as they are not hindered by the same limitations as text-based studies [16] and allow for analysis of the visual rhetoric constructed and utilized by political leaders [17]. Within the scope of visual research on political emotion display, two main schemes of analysis have shown to be especially valuable: the ethological approach [18] and the Facial Action Coding System (FACS) [19].
@@
Studies within the former largely focus on the analysis of political candidates’ non-verbal representations in news coverage, their displays during political campaigns and debates, as well as the impact of different forms of representation on viewers and success in political elections [20, 21, 22, 23, 24]. The Facial Action Coding System, on the other hand, allows for a more in-depth analysis of the (micro)expressive behavior of political leaders during speeches and public performances. Due to its complexity, Ekman’s classification system has been applied in a limited number of studies [25, 26, 27] but as the process of visual analysis gradually becomes automated [28] this research method is becoming more accessible to researchers from various fields of interest [29, 30].
@@
Our study expands on the existing framework of FACS research on political emotion display and applies a computer vision approach for detecting emotions from the facial expressions of political leaders. Because it is relatively new, this approach faces various methodological challenges which we aim to address by following the recommendations outlined by Barrett et al. [31], such as using big data techniques and automated detection in naturalistic settings, integrating traditional methods with machine learning, and sampling visual data across diverse contexts and cultures.
@@
Since the introduction of the FER dataset [32], different machine learning algorithms have been developed for solving the challenge of facial expression recognition, which essentially requires performing two tasks: (1) detection of facial expression, and (2) emotion recognition from the detected facial expressions. We develop a reproducible and principled computational workflow in Python for the analysis of video content of political leaders’ public performances. Our work relies on the existing Python library fer [33] which bundles a pre-trained convolutional neural network (CNN) Keras models for face detection [34, based on] and emotion recognition [35, 36].
@@ 2 Populism as performance
Defining populism represents a challenging task for researchers as there is barely any consensus on what such a definition should (or should not) include, and the concept itself becomes increasingly muddled as the term is used to describe a growing number of phenomena [37]. Most conceptualizations of populism can, however, be classified [38] into one of three main approaches: (1) the ideational approach [39], (2) the discursive approach [4, 40] and (3) the political-strategic approach [41]. For the purposes of this study, we adopt the discursive approach, which conceives populism primarily as a rhetorical style based on first-order principles pertaining to who should rule—the people, as “the only legitimate source of political and moral authority in a democracy”—but which says very little about the how [4]. Whereas the ideational approach claims populism is a “thin-centered” ideology [39] and emphasizes the interpretation of politics as “a Manichean struggle between a reified will of the people and a conspiring elite” [42] (Hawkins and Rovira Kaltwasser, 2018), the discursive approach highlights the ways in which this interpretation is framed and communicated as an anti-establishment discourse [43]. We find this to be the most comprehensible understanding of populism as it considers the operationalized elements of the populist communication logic [44] and examines the rhetorical components implicit in the ideological definition [45].
@@
From this perspective, the populist rhetoric has widely been characterized by the use of “dramatization, polarization, moralization, directness, ordinariness, colloquial and vulgar language” [44], and low-order appeals within which leaders “use a language that includes slang or folksy expressions and metaphors, are more demonstrative in their bodily or facial expressions as well as in their demeanor, and display more raw, culturally popular tastes” [40]. Because it represents a form of symbolic resistance to the more formalized and composed self-representation of mainstream pluralist leaders, i.e., the “establishment” and “political elite”, the populist performance  [^1] By “the populist performance” we mean the populist communication style as laid out by Bucy et al. [20]—that is, the verbal, tonal, and non-verbal display of a political leader which features three main communicative dimensions: simplification, emotionalization, and negativity. is subversive to the normative communication structures of political dialogue. As such, it appeals to “common sense” [20] and claims to give a voice to “the ordinary people” [46].
@@
This mode of communication has shown to be highly effective as a means of garnering media attention and cultivating a dedicated base of supporters—especially in the new media environment—and populist leaders seem to be overall more adept at capitalizing on it than other, non-populist political figures [9, 47]. Since digital media hinges on the attention economy [48] and thus tends to favor “dark participation” [49], populist leaders and political parties may fare well online due to the transgressive nature of their self-representation and the simplification, emotionalization, and negativity inherent in populist messaging [20, 44]. Social media is a particularly convenient tool for the proliferation of populist messages and imagery as it allows political actors to circumvent the mainstream media, reach large audiences, and communicate and interact directly with ‘the people’ [50, 51]. At the same time, online platforms allow for an instant audience response, as the “feedback loop between producer and consumer is much more intense and rapid than with television and radio” [52]. This form of ‘mediatized populism’ [53] allows for aspecific type of political performance, one which is largely unmitigated and relies heavily on portraying political leaders favorably to (online) audiences by framing them as expressive, dominant, and disruptive figures who are in tune with “the will of the people”.
@@
In recent years, much attention has been paid to the strategies of (affective) populist communication online, especially through analyses of verbal content shared by populist leaders and parties on Twitter and Facebook [e.g. 54, 55, 13, 14, 7]. Given the rise of visual communication in politics and the rapid developments of Web 2.0 practices [23, 49, 56], our primary aim was to expand the scope of research on the affective self-representation of populist leaders online by examining the characteristics of their non-verbal emotional communication. The facial expressions of political leaders are a particularly valuable area of research in this context, since digital media provide a breadth of material for analysis—well-lit, high-quality close-up shots of political leaders are now more prevalent than ever. We used this abundance of available data sources to our advantage and processed 220 YouTube videos uploaded to official accounts of political parties and their leaders, amounting to a total of 77 hours of video material. Our analysis examines potential differences in facial expressions between populist and non-populist political leaders during public performances, and more specifically, the characteristics of their negative emotional expressions and neutrality in emotional stance (i.e., absence of emotional expression).
@@ 3 Data & Method
3.1 Party selection
To create our sample, we first had to differentiate and categorize political parties (and, by extension, their leading figures) as either populist or pluralist. For this purpose, we used data from the Global Party Survey (GPS) [57], an expert survey from 2019 which offers an overview of key ideological and issue positions for 1043 parties from 163 countries worldwide. While expert surveys certainly have their limitations [58], we believe this database remains the most comprehensive comparative overview of political parties globally.
@@ 3.1 Party selection
Among other data, such as ideological framework and party size, the survey also offers an estimate of the degree to which existing parties employ populist rhetoric. In the questionnaire design, populist rhetoric is defined as “language which typically challenges the legitimacy of established political institutions and emphasizes that the will of the people should prevail” [57], whereas pluralist rhetoric “rejects these ideas, believing that elected leaders should govern, constrained by minority rights, bargaining and compromise, as well as checks and balances on executive power” [57]. The populist rhetoric scale (V8 in GPS dataset) is based on expert assessments of rhetorical style given these parameters. The variable Type_Populism categorizes this scale into four ordinal groups: strongly pluralist, moderately pluralist, moderately populist, and strongly populist. Although these categories may not be able to fully capture the nuances of populist rhetoric—its culturally specific manifestations, or what “the will of the people” may connote for people across different political contexts—they allow insight into the fundamental characteristics of party rhetoric and enable fair comparison between countries.
Given that a great concern in expert surveys is the possibility of experts’ backgrounds introducing cognitive bias into their assessments [59], it is also important to note that internal validity testing done within the GPS showed that the personal characteristics of experts (such as ideological stance, gender, and nationality) were not significant predictors when it comes to the placement of parties on the Pluralism-Populism scale [57]. [^2] For external validity measures, see robustness tests in Norris [57]. Additionally, to establish that differences in expression of emotion are indeed more of a matter of populist leaning rather than ideological preference, we use the “Type_Values” variable in the supplementary analysis to examine ideological differences in emotion expression (see the Supplement section E – Ideological differences).
@@
From the GPS dataset, fifteen countries with competitive popular elections were chosen: Argentina, Australia, Brazil, Croatia, France, Hungary, India, Italy, Japan, the Netherlands, Serbia, South Africa, Turkey, the United Kingdom, and the United States. [^3] The average number of experts per country in the selected dataset was approximately 27, with the highest number being in the United States (50) and the lowest in Serbia (6). While this number is higher than the average for the GPS as a whole—19 [57]—the variance between countries could potentially entail variance in overall quality of assessment. This sample, although limited, provides somewhat broad coverage of various cultures and political systems globally. Three parties with varying degrees of populist rhetoric—strongly populist, strongly pluralist, and either moderately populist or moderately pluralist—were selected from each country. We primarily focused on minor and major political parties as defined by the GPS variable categorizing parties based on their share of the vote in contests for the lower house of the national parliament. Fringe parties (less than 3% of the vote) were generally excluded from the sample since their audience reach and influence is considerably less than that of their larger counterparts competing at the same scale in their respective countries. Likewise, due to this criteria, only two parties from the US were included in the sample. Table from the Supplement D – Political leaders contains information about the selected parties and their leaders.
3.2 Video selection
@@
Once party selection was complete, we identified one leader per party who has held a prominent political position over the past decade (up to 2019, covered by the GPS), and found their official YouTube channel or their party’s official YouTube channel. From these channels, five videos were selected for each leader based on the following criteria:
@@
    View count. Videos that reached a larger audience were given priority over videos that did not perform comparatively as well on the platform. View count was used as a filtering mechanism to single out videos with large and potentially diverse audiences, because the most viewed videos on the official YouTube channels of major political parties usually refer to topics of wider public interest. This is partially due to the “popularity bias” on YouTube, meaning it has a tendency to recommend political content that aligns with the interests of the majority [60].
@@ 3.3 Video processing and analysis
Each video in the sample was processed by extracting 300 frames uniformly distributed over the length of the video, which were the basis for our main analysis. To improve the robustness of the overall approach, we also included an alternative frame processing approach in the supplementary material, based on processing every 50th frame of the video. Given that the majority of the YouTube videos have 24–30 frames per second, this is equivalent to selecting a frame every 1.67–2.00 seconds, which is a reasonable time window if we expect to capture potential changes in facial expression. The average number of frames per video extracted using this approach was 273.03, which makes the resulting time-series of emotion scores comparable in length with the first approach.
@@
For each selected frame of every video in the sample the following procedure is applied. First, face detection is applied to the extract frames using a three-stage cascaded CNN architecture with default parameters (scale factor: 0.709, minimum face size: 40 pixels, thresholds: [0.6, 0.7, 0.7] for the three detection stages) [34]. This approach is particularly suitable for detecting faces in videos because it performs simultaneous face detection and alignment, which is important as we cannot expect that political leaders will be looking straight at the camera in each extracted video frame. After face detection was applied to all frames, there were videos in which multiple faces were detected throughout the frame (when a second face appeared in >5% of frames). These videos were manually checked to confirm which of the detected faces corresponds to the selected political leader, and in some cases, specific frame ranges were specified to use only portions with consistent face detection.
@@
The second step is emotion detection within the verified detected face box. We relied on implementing the mini-Xception architecture of CNN [35] aimed at detecting 6 basic facial expressions of emotion, as well as a neutral expression. This model uses 4 residual depth-wise separable convolutions with input shape of 64x64 pixels (grayscale) and outputs 7 emotion scores. We initialized the detector with MTCNN enabled using the FER(mtcnn=True) constructor from the fer library. Similar to the face detection approach outlined above, this implementation is aimed at real-time emotion detection from video input for robotic applications. This model achieves 66% accuracy [see 35, p. 3 for more details] on the FER2013 dataset [32] which is a standard benchmark for emotion detection. This dataset consists of 30,000 small images (48x48 pixels) showing diverse facial expressions of different emotions.  [^4] FER2013 was constructed using Google Images API queries for queries for emotion-related keywords like “blissful” and“enraged”, combined with queries introducing diversity in gender, age and ethnicity represented in the images. Images were then labeled by human coders [32, pp. 3-4]. Unfortunately, no information is provided by the authors of the dataset regarding the number of human coders nor their background, or the labeling methodology applied to extract training labels for every image included in the FER2013 dataset. Benchmarks based on the FER2013 dataset should be taken with caution, especially in social science applications, because the dataset construction and validation documentation lacks crucial information about inter-coder reliability, coding procedures, and potential biases in the image selection process Each emotional expression has approximately 6000 images, except disgust, which is only present in 600 images. According to Goodfellow et al. [32], human classification accuracy is in the range of $65\%\pm 5\%$, but no additional information about the sample size, number of coders, or detection procedure is given.
@@
In order to provide additional information and to validate the output of the machine learning approach, we conducted a small image annotation study where 5 coders annotated a sample of 514 randomly selected frames extracted from 220 videos. Coders received basic training in emotion recognition and, due to limited resources, answered only two questions about every image: (1) Can you clearly see the face of a politician in this image? (Yes/No), (2) Is this face expressing positive emotion, negative emotion, or is it a neutral expression? (“Positive emotion”, “Negative emotion”, “Neutral expression”). The majority vote label was set as the “human-coded label” and compared with the output of the machine learning algorithm, whose label is defined based on the highest score out of the sum of positive emotions, negative emotions, or the neutrality score. A detailed description and the results of the data annotation study can be found in Supplement F – Image annotation.
@@
Figure 1: Example frames showing a high detected score of negative emotions (left: angry 0.78, disgust 0.00, fear 0.09, happy 0.00, sad 0.07, surprise 0.00, neutral 0.04) and a high detected score of neutral expression (right: angry 0.07, disgust 0.00, fear 0.04, happy 0.00, sad 0.02, surprise 0.08, neutral 0.79).
@@
3.4 Data processing pipeline
Our complete data processing pipeline consists of several sequential steps:
  1. 1.
    Video acquisition: Videos were downloaded from YouTube using the pytube library.
  2. 2.
    Frame extraction: Using the OpenCV (cv2) library, we extracted video metadata (fps, frame count, duration) and selected frames according to our sampling method.
  3. 3.
    Face detection: The MTCNN model implemented in the fer library detected and aligned faces in each frame.
  4. 4.
    Face verification: Videos with multiple detected faces were manually reviewed to identify the political leader.
  5. 5.
    Emotion recognition: The mini-Xception CNN model processed each detected face to output 7 emotion scores.
  6. 6.
    Data aggregation: For each video, we calculated mean scores for each emotion category and the composite negative emotion score.
  7. 7.
    Statistical analysis: We performed t-tests, ANOVA with post-hoc Tukey HSD tests, and calculated effect sizes to examine differences between populist and pluralist leaders.
All code for this pipeline is organized in a publicly available repository with a structured organization where raw data is processed through these steps, with intermediate and final results stored in appropriate directories. Videos requiring specific handling (e.g., cropping to specific time ranges) were processed using the moviepy library before entering the main pipeline.